# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.1 Lie dialgebra framework
$R_{\pm}=\frac{1}{2}(R\pm\text{id})$ (3.27)
which can be related to the projectors onto $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$ relative to direct sum decomposition (3.24) as
$R_{\pm}=\pm P_{\pm}\,.$ (3.28)
The linear map $R$ is related to the non-skew-symmetric $r$-matrix (1.13) underlying the cyclotomic Gaudin model, as one would naturally expect from the construction. More precisely, the expression (1.13) provides the kernel of the linear map (3.26) with respect to the bilinear pairing (3.17). To show this, we will use standard tensor product space notation as follows. Given any rational function $U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)$ in the parameters $\lambda$ and $\mu$ such that $\iota_{\mu}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}},\iota_{\lambda}\iota_{\mu}U_{{}_{\mathbf{\underline{12}}}}\in\widetilde{\mathfrak{g}}_{D}^{(0)}\,\widetilde{\otimes}\,\widetilde{\mathfrak{g}}_{D}^{(1)}$, where the first tensor factor in the formally completed tensor product is the loop algebra $\widetilde{\mathfrak{g}}_{D}^{(0)}$ with the loop parameter $\lambda$, and the second tensor factor is the space $\widetilde{\mathfrak{g}}_{D}^{(1)}$ with the loop parameter $\mu$, we define
${\langle\iota_{\mu}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}$ $=T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\left(\iota_{\mu_{r}}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\right)d\mu$
$\quad+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\left(\iota_{\mu_{r}}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\right)d\mu\,,$ (3.29)
@@
${\langle\iota_{\lambda}\iota_{\mu}U_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}$ $=T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\left(\iota_{\lambda}\iota_{\mu_{r}}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\right)d\mu$
$\quad+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\left(\iota_{\lambda}\iota_{\mu_{r}}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\right)d\mu\,,$ (3.30)
@@
$R_{+}(X)={\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}\quad\text{and}\quad R_{-}(X)={\langle\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}$ (3.31)
@@
$r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}\,.$ (3.32)
@@
In what follows, to treat the origin on the same footing as the points $\zeta_{r}$, $r\in\{1,\ldots,N\}$, it will be convenient to introduce the notation $\zeta_{0}=0$. We have
$\iota_{\lambda_{s}}\left(\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}\right)=\begin{dcases}\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{-km}(\lambda-\zeta_{s})^{m}\sigma_{{}_{\mathbf{\underline{2}}}}^{-k}C_{{}_{\mathbf{\underline{12}}}}}{(\mu-\omega^{-k}\zeta_{s})^{m+1}}&\text{for}\,\,s\in\{0,1,\ldots,N\}\\ -\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{k(m+1)}\mu^{m}\sigma_{{}_{\mathbf{\underline{2}}}}^{-k}C_{{}_{\mathbf{\underline{12}}}}}{\lambda^{m+1}}&\text{for}\,\,s=\infty\,,\\ \end{dcases}$ (3.33)
@@
$\iota_{\mu_{r}}\left(\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}\right)=\begin{dcases}-\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{k(m+1)}(\mu-\zeta_{r})^{m}\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{(\lambda-\omega^{k}\zeta_{r})^{m+1}}&\text{for}\,\,r\in\{0,1,\ldots,N\}\\ \frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{-km}\lambda^{m}\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu^{m+1}}&\text{for}\,\,r=\infty\,.\\ \end{dcases}$ (3.34)
@@
$W=\left(\sum_{n=0}^{\infty}W_{0}^{(n)}\lambda^{n},\sum_{n=0}^{\infty}W_{1}^{(n)}(\lambda-\zeta_{1})^{n},\ldots,\sum_{n=0}^{\infty}W_{N}^{(n)}(\lambda-\zeta_{N})^{n},\sum_{n=1}^{\infty}W_{\infty}^{(n)}\lambda_{\infty}^{n}\right)\in\widetilde{\mathfrak{g}}_{D+}^{(0)}\,,$ (3.35)
@@
$Z=\iota_{\lambda}\left(\sum_{n=0}^{N_{0}-1}\frac{Z_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\sigma^{k}Z_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Z_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right)\in\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}\,.$ (3.36)
Using the expansion (3.33), we find that
${\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},W_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=W\,.$ (3.37)
To evaluate ${\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}$, we note that for an arbitrary equivariant rational function $f\in\mathcal{F}_{D}^{\Gamma}$ and equivariant rational one-form $g\in\Omega_{D}^{\Gamma}$, we have the following relation:
$\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(g(\lambda)f(\lambda))d\lambda$ $=\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\sigma^{k}(g(\lambda))\sigma^{k}(f(\lambda)))d\lambda$
$=\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\omega^{k}g(\omega^{k}\lambda)f(\omega^{k}\lambda))d\lambda$
$=\operatorname{Res}_{\lambda_{r,k}=0}\operatorname{Tr}(g(\lambda)f(\lambda))d\lambda\,,$ (3.38)
for all $k\in\{0,\ldots,T-1\}$. This allows us to rewrite each entry of the tuple ${\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}$ as a sum over all the residues of a meromorphic one-form on $\mathbb{C}P^{1}$, and hence we deduce
${\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=0\,.$ (3.39)
@@
${\langle\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=W=P_{+}(X)=R_{+}(X)\,.$ (3.40)
Since $X\in\widetilde{\mathfrak{g}}_{D}^{(0)}$ was arbitrary, this establishes the first relation in (3.31). Similarly, using the expansion (3.34), we find ${\langle\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},W_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=0$ and ${\langle\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=-Z$ from which we conclude that
${\langle\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\rangle}_{{}_{\mathbf{\underline{2}}}}=-Z=-P_{-}(X)=R_{-}(X)\,.$ (3.41)
@@
$(\iota_{\mu}\iota_{\lambda}+\iota_{\lambda}\iota_{\mu})r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)\quad\text{and}\quad(\iota_{\mu}\iota_{\lambda}-\iota_{\lambda}\iota_{\mu})r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)$
@@
Recall that in Proposition 3.2 we identified the Lax matrix (1.16) of the cyclotomic Gaudin model, or rather its image under the embedding $\iota_{\lambda}$ in (3.18), as living in the subspace $V_{-}=\iota_{\lambda}\Omega_{D}^{\Gamma}$ of our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$. The next proposition establishes that this subspace $V_{-}\subset V$ is the orthogonal complement of $\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}\subset\widetilde{\mathfrak{g}}_{D}^{(0)}$, i.e., $\mathbb{g}_{-}\subset\mathbb{g}$, with respect to the nondegenerate bilinear pairing of Proposition 3.1. This completes the proof of the claim that the decomposition $V=V_{+}\dotplus V_{-}$ obtained in Proposition 3.2 is the one induced from the corresponding decomposition of the Lie algebra $\mathbb{g}=\mathbb{g}_{+}\dotplus\mathbb{g}_{-}$.
@@
An element $Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ lies in $\iota_{\lambda}\Omega_{D}^{\Gamma}$ if and only if $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$. Moreover, $Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ lies in $\widetilde{\mathfrak{g}}_{D+}^{(1)}$ if and only if $\langle Y,X\rangle=0$ for all $X\in\widetilde{\mathfrak{g}}_{D+}^{(0)}$.
@@
This is a particular case of the $\Gamma$-equivariant strong residue theorem [VY1, Appendix A]. We recall the proof here in the present setting for completeness. We start by proving that if $Y\in\iota_{\lambda}\Omega_{D}^{\Gamma}$, then $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$. Let us pick elements $f\in\mathcal{F}_{D}^{\Gamma}$ and $g\in\Omega_{D}^{\Gamma}$, and let $X=\iota_{\lambda}f$ and $Y=\iota_{\lambda}g$. Since $g(\lambda)f(\lambda)d\lambda$ is a meromorphic one-form on $\mathbb{C}P^{1}$, by the residue theorem we have
$\sum_{r=1}^{N}\sum_{k=0}^{T-1}\operatorname{Res}_{\lambda_{r,k}=0}\operatorname{Tr}(\iota_{\lambda_{r,k}}g(\lambda)\iota_{\lambda_{r,k}}f(\lambda))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda))d\lambda=0\,.$ (3.42)

### v2

@@ 3.1 Lie dialgebra framework
$R_{\pm}=\frac{1}{2}(R\pm\text{id}),$
which can be related to the projectors onto $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$ relative to direct sum decomposition (3.13) as
$R_{\pm}=\pm P_{\pm}.$
The linear map $R$ is related to the non-skew-symmetric $r$-matrix (1.8) underlying the cyclotomic Gaudin model, as one would naturally expect from the construction. More precisely, the expression (1.8) provides the kernel of the linear map (3.15) with respect to the bilinear pairing (3.10). To show this, we will use standard tensor product space notation as follows. Given any rational function $U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)$ in the parameters $\lambda$ and $\mu$ such that $\iota_{\mu}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}},\iota_{\lambda}\iota_{\mu}U_{{}_{\mathbf{\underline{12}}}}\in\widetilde{\mathfrak{g}}_{D}^{(0)}\,\widetilde{\otimes}\,\widetilde{\mathfrak{g}}_{D}^{(1)}$, where the first tensor factor in the formally completed tensor product is the loop algebra $\widetilde{\mathfrak{g}}_{D}^{(0)}$ with the loop parameter $\lambda$, and the second tensor factor is the space $\widetilde{\mathfrak{g}}_{D}^{(1)}$ with the loop parameter $\mu$, we introduce the following shorthand notations:
${\bigl{\langle}\iota_{\mu}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}$ ${}=T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu,$
@@
${\bigl{\langle}\iota_{\lambda}\iota_{\mu}U_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}$ ${}=T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\lambda}\iota_{\mu_{r}}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\lambda}\iota_{\mu_{r}}U_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)X_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
@@
$R_{+}(X)={\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}\qquad\text{and}\qquad R_{-}(X)={\bigl{\langle}\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}},$ (3.16)
@@
$r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}.$
@@
In what follows, to treat the origin on the same footing as the points $\zeta_{r}$, $r\in\{1,\dots,N\}$, it will be convenient to introduce the notation $\zeta_{0}=0$. We have
$\iota_{\lambda_{s}}\Biggl{(}\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}\Biggr{)}$
=$\qquad{}=\begin{dcases}\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{-km}(\lambda-\zeta_{s})^{m}\sigma_{{}_{\mathbf{\underline{2}}}}^{-k}C_{{}_{\mathbf{\underline{12}}}}}{\bigl{(}\mu-\omega^{-k}\zeta_{s}\bigr{)}^{m+1}}&\text{for}\ s\in\{0,1,\dots,N\},\\ -\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{k(m+1)}\mu^{m}\sigma_{{}_{\mathbf{\underline{2}}}}^{-k}C_{{}_{\mathbf{\underline{12}}}}}{\lambda^{m+1}}&\text{for}\ s=\infty,\\ \end{dcases}$ (3.17)
@@
$\iota_{\mu_{r}}\Biggl{(}\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu-\omega^{-k}\lambda}\Biggr{)}$
=$\qquad{}=\begin{dcases}-\frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{k(m+1)}(\mu-\zeta_{r})^{m}\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\bigl{(}\lambda-\omega^{k}\zeta_{r}\bigr{)}^{m+1}}&\text{for}\ r\in\{0,1,\dots,N\},\\ \frac{1}{T}\sum_{k=0}^{T-1}\sum_{m=0}^{\infty}\frac{\omega^{-km}\lambda^{m}\sigma_{{}_{\mathbf{\underline{1}}}}^{k}C_{{}_{\mathbf{\underline{12}}}}}{\mu^{m+1}}&\text{for}\ r=\infty.\\ \end{dcases}$ (3.18)
@@
$W=\Biggl{(}\sum_{n=0}^{\infty}W_{0}^{(n)}\lambda^{n},\sum_{n=0}^{\infty}W_{1}^{(n)}(\lambda-\zeta_{1})^{n},\dots,\sum_{n=0}^{\infty}W_{N}^{(n)}(\lambda-\zeta_{N})^{n},\sum_{n=1}^{\infty}W_{\infty}^{(n)}\lambda_{\infty}^{n}\Biggr{)}\in\widetilde{\mathfrak{g}}_{D+}^{(0)},$
@@
$Z=\iota_{\lambda}\Biggl{(}\sum_{n=0}^{N_{0}-1}\frac{Z_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\sigma^{k}Z_{r}^{(n)}}{\bigl{(}\omega^{-k}\lambda-\zeta_{r}\bigr{)}^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Z_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\Biggr{)}\in\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}.$
Using the expansion (3.17), we get
$T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)W_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)W_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad{}=\sum_{m=0}^{\infty}W_{s}^{(m)}(\lambda-\zeta_{s})^{m},\qquad\text{when}\quad s\in\{0,1,\dots,N\},$
and
$T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)W_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)W_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad{}=\sum_{m=1}^{\infty}W_{\infty}^{(m)}\lambda_{\infty}^{m},\qquad\text{when}\quad s=\infty.$
So, we find that
${\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},W_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=W.$ (3.19)
To evaluate ${\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}$, we note that for an arbitrary equivariant rational function ${f\in\mathcal{F}_{D^{\prime}}^{\Gamma}}$ and equivariant rational one-form $g\in\Omega_{D^{\prime}}^{\Gamma}$, we have the following relation:
$\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(g(\lambda)f(\lambda)){\rm d}\lambda$ $=\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\sigma^{k}(g(\lambda))\sigma^{k}(f(\lambda))\bigr{)}{\rm d}\lambda$
$=\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\omega^{k}g(\omega^{k}\lambda)f(\omega^{k}\lambda)\bigr{)}{\rm d}\lambda$
$=\operatorname{Res}_{\lambda_{r,k}=0}\operatorname{Tr}(g(\lambda)f(\lambda)){\rm d}\lambda$ (3.20)
for all $k\in\{0,\dots,T-1\}$. This allows us to rewrite each entry of the tuple ${\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}$ as
$T\sum_{r=1}^{N}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)Z_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)Z_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad{}=\sum_{r=1}^{N}\sum_{k=0}^{T-1}\operatorname{Res}_{\mu_{r,k}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)Z_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu$
$\qquad\quad{}+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\mu_{r}=0}\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}\bigl{(}\iota_{\mu_{r}}\iota_{\lambda_{s}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)Z_{r_{\mathbf{\underline{2}}}}(\mu_{r})\bigr{)}{\rm d}\mu,\qquad s\in S,$
which is a sum over all the residues of a meromorphic one-form on $\mathbb{C}P^{1}$. Hence, we deduce
${\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=0.$ (3.21)
@@
${\bigl{\langle}\iota_{\mu}\iota_{\lambda}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=W=P_{+}(X)=R_{+}(X).$
Since $X\in\widetilde{\mathfrak{g}}_{D}^{(0)}$ was arbitrary, this establishes the first relation in (3.16). Similarly, using the expansion (3.18), we find ${\bigl{\langle}\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},W_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=0$ and ${\bigl{\langle}\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},Z_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=-Z$ from which we conclude that
${\bigl{\langle}\iota_{\lambda}\iota_{\mu}r_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{2}}}}\bigr{\rangle}}_{{}_{\mathbf{\underline{2}}}}=-Z=-P_{-}(X)=R_{-}(X).$
@@
$(\iota_{\mu}\iota_{\lambda}+\iota_{\lambda}\iota_{\mu})r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)\qquad\text{and}\qquad(\iota_{\mu}\iota_{\lambda}-\iota_{\lambda}\iota_{\mu}),r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),$
@@
Recall that in Proposition 3.2 we identified the Lax matrix (1.9) of the cyclotomic Gaudin model, or rather its image under the embedding $\iota_{\lambda}$ in (3.11), as living in the subspace ${V_{-}=\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}}$ of our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$. The next proposition establishes that this subspace $V_{-}\subset V$ is the orthogonal complement of $\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}\subset\widetilde{\mathfrak{g}}_{D}^{(0)}$, i.e., $\mathbb{g}_{-}\subset\mathbb{g}$, with respect to the nondegenerate bilinear pairing of Proposition 3.1. This completes the proof of the claim that the decomposition $V=V_{+}\dotplus V_{-}$ obtained in Proposition 3.2 is the one induced from the corresponding decomposition of the Lie algebra $\mathbb{g}=\mathbb{g}_{+}\dotplus\mathbb{g}_{-}$.
@@
An element $Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ lies in $\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$ if and only if $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$. Moreover, $Y\!\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ lies in $\widetilde{\mathfrak{g}}_{D+}^{(1)}$ if and only if $\langle Y,X\rangle=0$ for all $X\!\in\widetilde{\mathfrak{g}}_{D+}^{(0)}$.
@@
This is a particular case of the $\Gamma$-equivariant strong residue theorem [41, Appendix A]. We recall the proof here in the present setting for completeness. We start by proving that if $Y\in\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$, then $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$. Let us pick elements $f\in\mathcal{F}_{D^{\prime}}^{\Gamma}$ and $g\in\Omega_{D^{\prime}}^{\Gamma}$, and let $X=\iota_{\lambda}f$ and $Y=\iota_{\lambda}g$. Since $g(\lambda)f(\lambda){\rm d}\lambda$ is a meromorphic one-form on $\mathbb{C}P^{1}$, by the residue theorem we have
$\sum_{r=1}^{N}\sum_{k=0}^{T-1}\operatorname{Res}_{\lambda_{r,k}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r,k}}g(\lambda)\iota_{\lambda_{r,k}}f(\lambda)\bigr{)}{\rm d}\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda=0.$